OpenStudy (anonymous):

What is 6 over x^3-64 minus 4 over x^2-16???

OpenStudy (anonymous):

\[\frac{6}{x^2-64}=\frac{6}{(x-4)(x^2+4x+16)} \] \[\frac{4}{x^2-16}=\frac{4}{(x-4)(x+4)}\] is a start

OpenStudy (anonymous):

typo there, first line should be \[\frac{6}{x^3-64}=\frac{6}{(x-4)(x^2+4x+16)}\]

OpenStudy (anonymous):

so least common multiple of the denominators is \[( x+4)(x-4)(x^2+4x+16)\]

OpenStudy (anonymous):

How would i write it all out? And get the answer?

OpenStudy (vishal_kothari):

can u draw the question?

OpenStudy (anonymous):

Yes hold on

OpenStudy (anonymous):